# MedMamba: Vision Mamba for Medical Image Classification

### v2

@@ Abstract
Medical image classification is a very fundamental and crucial task in the field of computer vision. These years, CNN-based and Transformer-based models have been widely used to classify various medical images. Unfortunately, The limitation of CNNs in long-range modeling capabilities prevents them from effectively extracting features in medical images, while Transformers are hampered by their quadratic computational complexity. Recent research has shown that the state space model (SSM) represented by Mamba can efficiently model long-range interactions while maintaining linear computational complexity. Inspired by this, we propose Vision Mamba for medical image classification (MedMamba). More specifically, we introduce a novel Conv-SSM module. Conv-SSM combines the local feature extraction ability of convolutional layers with the ability of SSM to capture long-range dependency, thereby modeling medical images with different modalities. To demonstrate the potential of MedMamba, we conducted extensive experiments using 14 publicly available medical datasets with different imaging techniques and two private datasets built by ourselves. Extensive experimental results demonstrate that the proposed MedMamba performs well in detecting lesions in various medical images. To the best of our knowledge, this is the first Vision Mamba tailored for medical image classification. The purpose of this work is to establish a new baseline for medical image classification tasks and provide valuable insights for the future development of more efficient and effective SSM-based artificial intelligence algorithms and application systems in the medical. Source code has been available at https://github.com/YubiaoYue/MedMamba.
@@ 1 Introduction
Influenced by the success of VMamba [22] in natural image classification tasks, we propose MedMamba for the first time, a simple model based on modern SSM, aiming to demonstrate the potential of Mamba in various medical image classification tasks. Specifically, the core of MedMamba is composed of a module called Conv-SSM. Unlike natural images, there is a high degree of similarity between various medical images, so the model is required to be able to effectively extract both fine-grained features and coarse-grained features. Based on this basic fact, Conv-SSM uses both the local feature extraction ability of the convolutional layer and the long-range modeling ability of SSM to achieve the effective extraction of coarse and fine-grained features in medical images and accurate image classification. Comprehensive experiments were conducted on five medical datasets acquired by different imaging techniques to demonstrate the potential of MedMamba modules in medical image classification. Specifically, we conduct extensive experiments on three public datasets and two private datasets, and the results show that MedMamba can achieve competitive performance.
@@ 2.2 MedMamba
As shown in Figure 1, the overall architecture of MedMamba is shown. Specifically, the structure of MedMamba is very simple, including only the patch embedding layer, Conv-SSM Block and patch merging layer.
In the MedMamba, a Patch Embedding layer firstly divides the input image $x\in\mathcal{R}^{H\times W\times 3}$ into non-overlapping patches of size $4\times 4$, thereby mapping the dimensions of the image to $C$ (Here, the default values of $C$ is set to 96), resulting in an embedded image $x^{\prime}\in\mathcal{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$. Subsequently, The $x^{\prime}$ is normalized by using a Layer Normalization [25] before feeding it into the backbone of MedMamba. The backbone is composed of four stages. It should be noted that patch merging layers are used after the output of the first three stages to reduce the height and width of input features, while increasing the number of channels. We use [2, 2, 2, 2] Conv-SSM blocks in four stages, with each stage having [C, 2C, 4C, 8C] channels.
@@
2.3 Conv-SSM Block
Conv-SSM Block is the core module of MedMamba, as shown in Figure 1. Overall, Conv-SSM is a dual-branch module. Firstly, the channel splitting operation is used to split the module input into two sub inputs with the same size. Next, the two sub inputs are fed into two branch (i.e., Conv-branch and SSM-branch)modules, respectively. In the Conv-branch, we simply use continuous convolutional layers to model the local features of the input. In the SSM-branch, we first utilize the layer Normalization layer to process the input. It is worth noting that the input is divided into two branches after passing through Layer Normalization. In the first branch of SSM, the input passes through a linear layer and an activation function. In the second branch of SSM-branch, the input is fed into the 2D Selective Scan module (SS2D) [22] for further feature extraction after passing through a linear layer, depthwise separable convolution, and activation function. Then, Layer Normalization is used to normalize the features and multiply them element by element with the output of the first branch, merging the two branches. Finally, a linear layer is used for feature mixing and added with residual connections to obtain the output of Conv-SSM Block. It is worth noting that the activation function in the SSM-branch defaults to SiLU[26]. At the end of the network, we merge the outputs of two branches along the channel dimension of the feature map and use a convolutional layer with a kernel size of 1 to fuse the feature map, thereby facilitating information interaction between channels.
@@ 3.4 Main results and findings
Dataset Model Fold-1(%)normal-$\uparrow$ Fold-2(%)normal-$\uparrow$ Fold-3(%)normal-$\uparrow$ Fold-4(%)normal-$\uparrow$ Fold-5(%)normal-$\uparrow$ AOA(%)normal-$\uparrow$
Cervical-US ResNet50[35] 80.66 80.75 80.64 80.80 81.43 80.86(±0.29)
ConvNeXt-B [36] 80.78 81.68 81.61 80.64 81.10 81.16(±0.42)
ViT-B [17] 82.57 79.76 78.73 76.96 77.63 79.13(±1.97)
Swin-T [37] 80.16 79.62 80.93 80.10 80.58 80.28(±0.44)
MedMamba 85.38 83.90 83.60 84.49 84.94 84.46(±0.65)
Fetal-US ResNet50[35] 89.87 89.27 89.10 88.70 89.23 89.23(±0.38)
ConvNeXt-B [36] 89.02 88.66 88.62 88.58 89.63 88.89(±0.40)
ViT-B [17] 90.39 90.15 89.10 91.81 88.22 89.93(±1.22)
Swin-T [37] 90.56 90.64 89.87 90.44 89.47 90.20(±0.45)
MedMamba 93.10 92.82 92.37 93.18 93.77 93.05(±0.46)
CPN-Xray MobileNetV3-L[38] 93.62 93.71 93.02 93.70 94.33 93.67(±0.42)
EfficientNetV2-M [39] 95.40 94.91 95.55 95.30 95.57 95.35(±0.24)
ViT-B [17] 94.29 94.35 94.09 94.76 94.41 94.38(±0.22)
Swin-T [37] 95.60 95.89 96.08 95.60 95.99 95.83(±0.20)
MedMamba 96.46 97.13 96.27 96.75 96.67 96.66(±0.29)
Kavsir ResNet50[35] 73.91 74.62 74.94 73.94 74.79 74.44(±0.43)
ConvNeXt-B [36] 74.63 74.23 74.02 73.99 73.59 74.09(±0.34)
ViT-B [17] 73.86 73.99 73.60 73.65 74.00 73.82(±0.17)
Swin-T [37] 74.71 75.04 75.75 74.66 74.49 74.93(±0.45)
MedMamba 74.25 74.63 74.63 73.38 74.50 74.28(±0.47)
Otoscopy2024 ResNet50 [35] 87.57 88.74 88.59 86.62 87.04 87.71(±0.84)
ConvNeXt-B [36] 90.94 89.64 87.88 88.09 88.15 88.94(±1.18)
Vit-B [17] 89.37 92.14 91.13 91.77 90.28 90.94(±1.00)
Swin-T [37] 90.10 89.66 88.87 89.90 90.52 89.81(±0.55)
MedMamba 91.36 90.55 90.05 89.60 90.52 90.41(±0.58)

### v3

@@ Abstract
Medical image classification is a very fundamental and crucial task in the field of computer vision. These years, CNN-based and Transformer-based models have been widely used to classify various medical images. Unfortunately, The limitation of CNNs in long-range modeling capabilities prevents them from effectively extracting features in medical images, while Transformers are hampered by their quadratic computational complexity. Recent research has shown that the state space model (SSM) represented by Mamba can efficiently model long-range interactions while maintaining linear computational complexity. Inspired by this, we propose Vision Mamba for medical image classification (MedMamba). More specifically, we introduce a novel SS-Conv-SSM module. SS-Conv-SSM combines the local feature extraction ability of convolutional layers with the ability of SSM to capture long-range dependency, thereby modeling medical images with different modalities. To demonstrate the potential of MedMamba, we conducted extensive experiments using 14 publicly available medical datasets with different imaging techniques and two private datasets built by ourselves. Extensive experimental results demonstrate that the proposed MedMamba performs well in detecting lesions in various medical images. To the best of our knowledge, this is the first Vision Mamba tailored for medical image classification. The purpose of this work is to establish a new baseline for medical image classification tasks and provide valuable insights for the future development of more efficient and effective SSM-based artificial intelligence algorithms and application systems in the medical. Source code has been available at https://github.com/YubiaoYue/MedMamba.
@@ 1 Introduction
Influenced by the success of VMamba [22] in natural image classification tasks, we propose MedMamba for the first time, a simple model based on modern SSM, aiming to demonstrate the potential of Mamba in various medical image classification tasks. Specifically, the core of MedMamba is composed of a module called SS-Conv-SSM. Unlike natural images, there is a high degree of similarity between various medical images, so the model is required to be able to effectively extract both fine-grained features and coarse-grained features. Based on this basic fact, SS-Conv-SSM uses both the local feature extraction ability of the convolutional layer and the long-range modeling ability of SSM to achieve the effective extraction of coarse and fine-grained features in medical images and accurate image classification. Comprehensive experiments were conducted on five medical datasets acquired by different imaging techniques to demonstrate the potential of MedMamba modules in medical image classification. Specifically, we conduct extensive experiments on three public datasets and two private datasets, and the results show that MedMamba can achieve competitive performance.
@@ 2.2 MedMamba
As shown in Figure 1, the overall architecture of MedMamba is shown. Specifically, the structure of MedMamba is very simple, including only the patch embedding layer, SS-Conv-SSM Block and patch merging layer.
In the MedMamba, a Patch Embedding layer firstly divides the input image $x\in\mathcal{R}^{H\times W\times 3}$ into non-overlapping patches of size $4\times 4$, thereby mapping the dimensions of the image to $C$ (Here, the default values of $C$ is set to 96), resulting in an embedded image $x^{\prime}\in\mathcal{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$. Subsequently, The $x^{\prime}$ is normalized by using a Layer Normalization [25] before feeding it into the backbone of MedMamba. The backbone is composed of four stages. It should be noted that patch merging layers are used after the output of the first three stages to reduce the height and width of input features, while increasing the number of channels. We use [2, 2, 4, 2] SS-Conv-SSM blocks in four stages, with each stage having [C, 2C, 4C, 8C] channels.
@@
2.2.1 2D-selective-scan
Figure 2: Illustration of the 2D-Selective-Scan on a feature map. We commence by scanning an image using scan expanding. The four resulting features are then individually processed through the S6 block, and the four output features are merged by scan merging into the final 2D feature map.
We first describe the 2D-selective-scan (SS2D) proposed by the VMamba [22], which is the core of SS-Conv-SSM Block. The SS2D consists of three components: a scan expanding operation, an S6 block, and a scan merging operation. As shown in Figure 2, the scan expanding operation unfolds the input image along four different directions (top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right) into sequences. These sequences are then processed by the S6 block for feature extraction, ensuring that information from various directions is thoroughly scanned, thus capturing diverse features. Subsequently, as illustrated in Figure 2, the scan merging operation sums and merges the sequences from the four directions, restoring the output image to the same size as the input. The S6 block, derived from Mamba [20], introduces a selective mechanism on top of S4 [26] by adjusting the SSM’s parameters based on the input. This enables the model to distinguish and retain pertinent information while filtering out the irrelevant. The pseudo-code for the S6 block is presented in Algorithm 1.
Algorithm 1 Pseudo-code for S6 block in SS2D [22, 24, 20]
0:  $x$, the feature with shape [B, L, D] (batch size, token length, dimension)
0:  $\mathbf{A}$, the nn.Parameter; $\mathbf{D}$, the nn.Parameter
0:  Linear(.), the linear projection layer
0:  $y$, the feature with shape [B, L, D]
1:  $\mathbf{\Delta},\mathbf{B},\mathbf{C}$ = Linear($x$), Linear($x$), Linear($x$)
2:  $\overline{\mathbf{A}}=\textup{exp}(\mathbf{\Delta}\mathbf{A})$
3:  $\overline{\mathbf{B}}=(\mathbf{\Delta}\mathbf{A})^{-1}(\textup{exp}(\mathbf{\Delta}\mathbf{A})-\mathbf{I})\cdot\mathbf{\Delta}\mathbf{B}$
4:  $h_{t}=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t}$
5:  $y_{t}=\mathbf{C}h_{t}+\mathbf{D}x_{t}$
6:  $y=[y_{1},y_{2},\cdots,y_{t},\cdots,y_{L}]$
7:  return  $y$
2.2.2 SS-Conv-SSM Block
SS-Conv-SSM Block is the core module of MedMamba, as shown in Figure 1. Overall, SS-Conv-SSM is a straightforward dual-branch module. There aren’t many fancy tricks in SS-Conv-SSM Block. Firstly, the channel splitting operation is used to split the module input into two sub inputs with the same size. Next, the two sub inputs are fed into two branch (i.e., Conv-branch and SSM-branch)modules, respectively. In the Conv-branch, we simply use continuous convolutional layers to model the local features of the input. In the SSM-branch, we first utilize the layer Normalization layer to process the input. It is worth noting that the input is divided into two branches after passing through Layer Normalization. In the first branch of SSM, the input passes through a linear layer and an activation function. In the second branch of SSM-branch, the input is fed into the 2D Selective Scan module (SS2D) [22] for further feature extraction after passing through a linear layer, depthwise separable convolution, and activation function. Then, Layer Normalization is used to normalize the features and multiply them element by element with the output of the first branch, merging the two branches. Finally, a linear layer is used for feature mixing, and obtain the output of SSM-branch. At the end of the SS-Conv-SSM block, we merge the outputs of two branches along the channel dimension of the feature map and use a lightweight operation without parameters, i.e.,channel shuffle operation to facilitate information interaction between channels of two sub inputs. It is worth noting that the activation functions in the SSM-branch and Conv-branch default to SiLU and ReLU, respectively[27, 28].
@@ 3.1 Datasets
PAD-UFES-20. The PAD-UFES-20 dataset was collected along with the Dermatological and Surgical Assistance Program (in Portuguese: Programa de Assistência Dermatológica e Cirurgica - PAD) at the Federal University of Espírito Santo (UFES-Brazil), which is a nonprofit program that provides free skin lesion treatment, in particular, to low-income people who cannot afford private treatment. The dataset consists of 2,298 samples of six different types of skin lesions. Each sample consists of a clinical image and up to 22 clinical features including the patient’s age, skin lesion location, Fitzpatrick skin type, and skin lesion diameter. The skin lesions are: Basal Cell Carcinoma (BCC), Squamous Cell Carcinoma (SCC), Actinic Keratosis (ACK), Seborrheic Keratosis (SEK), Bowen’s disease (BOD), Melanoma (MEL), and Nevus (NEV). As the Bowen’s disease is considered SCC in situ, we clustered them together, which results in six skin lesions in the dataset, three skin cancers (BCC, MEL, and SCC) and three skin disease (ACK, NEV, and SEK). The images present in the dataset have different sizes because they are collected using different smartphone devices. All images are available in .png format. In total, there are 1,373 patients, 1,641 skin lesions, and 2,298 images present in the dataset. Each image/sample has a reference to the patient and the skin lesion in the metadata (dataset url: https://data.mendeley.com/datasets/zr7vgbcyr2/1).
@@ 3.2 Evaluation Metrics
In addition, due to the characteristics of medical images, AUC (Area under the receiver operating characteristic curve) was also used to further measure model performance.
@@ 3.4 Main results and findings
Table 2: The performance of MedMamba on various datasets and comparison with reference models. (OA represents the overall accuracy.)
[…]
PAD-UFES-20 ResNet50[37] 45.71 42.40 89.78 42.77 56.62 0.7626
ConvNeXt-B [38] 33.80 34.47 88.97 33.36 54.73 0.7613
ViT-B [17] 32.04 33.21 88.05 32.20 50.36 0.7291
Swin-T [39] 49.35 43.10 90.83 42.87 62.15 0.8396
MedMamba 38.43 36.94 89.90 35.80 58.80 0.8070
Cervical-US ResNet50[37] 76.97 76.80 94.49 76.75 84.35 0.9433
ConvNeXt-B [38] 71.11 67.55 92.70 68.90 80.01 0.8875
ViT-B [17] 59.15 54.59 90.05 54.86 73.91 0.8412
Swin-T [39] 73.24 60.06 91.29 62.00 77.75 0.8804
MedMamba 82.67 73.83 94.68 76.32 85.62 0.9524
Fetal-US ResNet50[37] 91.48 92.29 98.53 91.85 93.08 0.9901
ConvNeXt-B [38] 84.62 81.40 97.21 82.69 87.27 0.9738
ViT-B [17] 77.95 77.18 96.51 77.50 83.66 0.9625
Swin-T [39] 87.39 87.02 97.95 87.10 90.34 0.9867
MedMamba 92.15 93.89 98.73 92.97 93.97 0.9931
CPN-Xray ResNet50[37] 97.57 97.55 98.74 97.56 97.51 0.9970
ConvNeXt-B [38] 93.82 93.84 96.85 93.83 93.75 0.9867
ViT-B [17] 92.09 91.98 95.87 91.97 91.83 0.9807
Swin-T [39] 95.64 95.61 97.73 95.62 95.53 0.9911
MedMamba 97.21 97.17 98.54 97.19 97.12 0.9953
Kavsir ResNet50[37] 77.99 78.08 96.86 77.82 78.08 0.9729
ConvNeXt-B [38] 74.09 74.24 96.32 73.87 74.25 0.9660
ViT-B [17] 74.46 73.25 96.17 72.40 73.25 0.9657
Swin-T [39] 73.34 73.16 96.16 72.84 73.16 0.9630
MedMamba 78.74 78.83 96.97 78.56 78.83 0.9731
Otoscopy2024 ResNet50 [37] 86.06 81.00 98.47 83.23 88.65 0.9847
ConvNeXt-B [38] 79.32 75.92 97.99 77.39 84.92 0.9782
Vit-B [17] 78.36 74.16 87.81 75.88 83.44 0.9739
Swin-T [39] 83.45 77.44 98.25 80.05 86.91 0.9821
MedMamba 86.00 84.44 98.59 85.15 89.45 0.9889